• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Higinbotham

    Log entry time 16:04:23 on August26,2009

    Entry number 286136

    keyword=Shift Summary

    Starting shift 5 uA thin Ta (4341) right arm only. Left arm HV's are off.

    Switching from thin to thick Ta.

    30 uA beam

    Controlled access at 10am, radcon cleared at 10:35am, Ramesh and Dustin enter hall.

    14:00 Everyone is out of the Hall. At the same time MCC is ready for tuning.

    15:30 ERS went down. It came back in time to save the target, but we lost 4 out of 6 spectrometer magnets.

    16:00 Controlled access to reset magnets.